OpenStudy (anonymous):

Multiply. (x – 4)(x2 + 5x + 6)

OpenStudy (anonymous):

x(x2 + 5x + 6) -4(x2 + 5x + 6) = x3 + 5x2 + 6x -4x2 - 20x -24 = x3 +x2 -14x - 24